# DIAS: A Comprehensive Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

### 12 Dec 2023 v2

@@ I Introduction
Cerebrovascular diseases significantly contribute to global mortality and long-term disability [1]. In recent decades, there has been a notable decline in the mortality rate associated with intracranial artery (IA) diseases, such as intracranial arterial stenosis (ICAS), middle cerebral artery occlusion (MCAO), and intracranial aneurysm, largely attributed to rapid advancements in medical and imaging technology [2]. Digital Subtraction Angiography (DSA), an imaging modality, visualizes the flow of a contrast agent through vessels by capturing a sequence of images over a duration typically spanning 3 to 15 seconds, with a sampling rate between 3 to 5 frames per second [3]. It removes bony structures by subtracting the pre-contrast image, thus preserving the vessels filled with contrast agent. Owing to its inherently superior spatial and temporal resolution, DSA can accurately elucidate lesion details in instances where computed tomography angiography (CTA) and magnetic resonance angiography (MRA) may not offer a definitive diagnosis [2]. Thus, DSA is widely recognized as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
@@
Conventionally, neuroradiologists undertake the visual analysis of DSA for diagnostic purposes, clinical decision-making, and postoperative assessment, which is inherently time-consuming, labor-intensive, and subjective. Automated segmentation of vessels DSA harbors the potential to mitigate this challenge by delineating vascular morphology. Currently, DSA-based cerebrovascular image segmentation finds utility in various downstream clinical applications, including automated thrombolysis in cerebral infarction scoring [5], 3D reconstruction of vessel [6], and image guidance of interventional surgery [7]. Despite the undeniable significance of cerebrovascular segmentation in DSA, it remains an under-researched area. One possible reason is the current absence of a DSA dataset tailored specifically for the formulation and evaluation of cerebrovascular segmentation techniques in the medical imaging community. However, given the intrinsic attributes of its imaging modality, the task of creating a brain DSA vascular segmentation dataset is immensely challenging. Primarily, the DSA sequence presents a projection of intravascular contrast media from a consistent viewing angle, thereby visualizing only a fraction of the IA within the contrast agent in any given image, as illustrated in Fig. 1. In the process of annotating vessels within DSA, neuroradiologists are required to conduct a detailed examination of each image frame, which is both laborious and expensive [8]. Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures in imagery, diminished contrast, and multifaceted background disturbances, accentuates the difficulties associated with annotation.
@@
Currently, several segmentation approaches for DSA cerebral vessels have emerged [9, 10, 8]. Yet, all these efforts are anchored to proprietary datasets, absent a consistent evaluation benchmark. Furthermore, most of these menthods [9, 10] are limited to single-frame image segmentation, overlooking the rich temporal characteristics inherent in DSA sequences. Segmentation methods based on single-frame DSA may not comprehensively capture the vascular intricacies of IA, potentially limiting their efficacy in diagnosing and treating cerebrovascular diseases [8]. In DSA sequences, vascular segmentation is characterized as a dimensionality-reduction task, taking a 2D+time image input to yield a 2D mask output. This task deviates from traditional 2D and 3D vascular segmentation meathods [11, 12, 13]. The pivotal challenge lies in accurately capturing the spatiotemporal information present within the angiography for precise vessel segmentation.
@@
Figure 1: A representative sample of intracranial vessels in sagittal view from DSA, where t to t+5 are consecutive frames of the arterial phase.
@@
In this work, we construct the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons. Given the intricate and labor-intensive nature of vessel annotation within DSA sequences, we’ve provided two types of weak annotations for training samples, aiming to facilitate studies on weakly-supervised IA segmentation. For the DIAS, we have devised three benchmarks tailored for fully-supervised segmentation (FSS), weakly-supervised segmentation (WSS), and semi-supervised segmentation (SSS): the vessel sequence segmentation network, the scribble learning-based segmentation framework and the random patch-based self-training framework. We have carried out comprehensive experiments using the DIAS, establishing robust baselines that will serve as a foundation for future algorithmic comparisons on the DIAS dataset. We hope that the DIAS dataset could facilitate open research in cerebral vessel segmentation.
@@ II-A Vessel Segmentation Datasets
Vessel segmentation is a popular area of research within computer-aided medicine, which has witnessed remarkable progress in recent years, owing to technological advancements and the availability of high-quality annotated datasets. We have summarized the several publicly available vessel segmentation datasets in Table I. A representative example is retinal vessel segmentation, which has many publicly available datasets due to its easy availability, such as colour fundus datasets (DRIVE [14], CHASE_DB1 [15], STARE [16], HRF [17]) and OCTA datasets (ROSE1 and ROSE2 [18]). These datasets consist of 20-117 2D images with spatial resolution ranging from 304$\times$304 to 3504$\times$2336. However, other vessel segmentation datasets are relatively scarce [8]. CHUAC [19] and DCA1 [20] are coronary angiography datasets with resolutions of 189$\times$189 and 300$\times$300, respectively. ASOCA [21] and VESSEL12 [22] are the 3D vessel segmentation datasets of coronary artery CTA and lung CT from the challenges, in which only training set data is available.
The availability of such datasets has led to a surge of outstanding works in the field of segmentation, with researchers striving to develop algorithms that outperform existing methods. So far, there have been no publicly available DSA datasets with vessel annotation in the medical imaging community, and only a few studies conducted with in-house datasets. DCVessel [9] comprises 30 DSA arterial-phase images at a resolution of 700$\times$700, while HEART [23] comprises 1092 DSA coronary images at 256$\times$256 resolution. Su et al. [8] manually segmented arteries and veins on 97 DSA series, covering both anterior-posterior and lateral views, each with a resolution of 1024$\times$1024 pixels. In this work, we have released DIAS, an IA segmentation dataset derived from DSA sequences. It comprises 120 sequences (753 frames), each with a resolution of 800$\times$800 pixels, and the lengths of these sequences range from 4 to 14. The DIAS can serve as a reliable benchmark for the development of algorithms on IA segmentation from DSA sequences. We hope that this dataset will intrigue researchers to explore this field further.
@@ II-B Vessel Segmentation Methods
In the past few years, deep learning has demonstrated great potential in vessel segmentation, surpassing traditional methods, such as hand-crafted features [24], filtering-based models [25], and statistical models [26]. Deep learning-based methods have become primary approaches for vessel segmentation, with UNet [27] serving a crucial role as a milestone network architecture, giving rise to numerous variant networks. These networks utilize an encoder-decoder architecture with popular components, such as multi-scale design [13, 9, 28], deep supervision [29, 9, 12], attention mechanisms [30, 13, 31], atrous convolution [32, 29, 12, 9], and improved skip connections [9, 12]. Moreover, several novel neural networks [33, 34] have also succeeded in vessel segmentation by leveraging generative adversarial networks (GAN) or graph neural networks (GNN). Some studies focused on vessel segmentation beyond network architecture, such as optimized training strategies [35, 36] and modifying loss functions [37].
@@
However, there have been few studies on DSA vessel segmentation. Meng et al. [9] proposed a CNN-based multiscale dense CNN segmentation framework, which includes a multiscale atrous convolution module, an improved dense block, and redesigned skip connections, for IA segmentation in single frame DSA. Zhang et al. [23] proposed a boundary enhancement and feature denoising module to improve the extraction of boundary information in DSA coronary segmentation. These two approaches are tailored for vessel segmentation in individual DSA frames. For vessel sequence, Hao et al. proposed the sequential vessel segmentation network (SVS-Net) for vessel segmentation in X-ray coronary angiography, which incorporates 3D convolutional layers to extract rich temporal-spatial features and employs channel attention blocks to learn discriminative features [38]. Su et al. proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences [8]. While both methods achieve vessel segmentation within sequences, the former is not adaptable to variable sequences, and the latter is challenging to extend to other models.
@@
Fully annotating vessel datasets is both time-consuming and costly. Consequently, annotation-efficient segmentation methods based on weakly-supervised and semi-supervised have attracted significant attention from researchers. WSL trains models with sparse annotations instead of full ones. The common methods involves online labeling [39] via consistency checks with the mean teacher or using pseudo-labels supervision [40] derived from model predictions. SSS aims to explore tremendous unlabeled data with supervision from limited labeled data. Recently, self-training methods [41, 42] have dominated this field. Furthermore, methods employing consistency regularization strategies [41, 43, 44] improve the generalization ability by encouraging high similarity in predictions from two perturbed networks for the same input image. To our knowledge, there is no existing literature on semi-supervised or weakly-supervised vessel segmentation in DSA sequences. In this regard, we propose a comprehensive benchmark involving FSS, WSS, and SSS for IA segmentation based on the constructed DSA-sequences dataset.
Figure 2: Full annotated process of DIAS.
Figure 3: Full annotated representative samples of DIAS at coronal view and sagittal view.
@@ III-A Dataset Description
The DIAS dataset was derived from intraoperative DSA images obtained during ischemic stroke treatments, including ICAS and MCAO, at Beijing Tiantan Hospital. The ethical committee at Beijing Tiantan Hospital authorized the use of all the data presented in this study. All images were retrospectively collected between January 2019 and December 2021 and then anonymized to protect the patients’ privacy by removing all identifiable clinical treatment details. The original DSA sequences were captured by the angiography machine during the actual endovascular therapy process and stored as DICOM files. It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view, lateral view or both. In this study, we retained 120 sequences after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates. The number of frames varies between sequences (6—46 frames), but the frame rate is fixed (4 FPS). The sequences were then extracted into individual images with a resolution of 800$\times$800 pixels using RadiAnt. We retained all the arterial-phase frames and removed the pre-contrast, capillary-phase, and venous-phase frames under the guidance of a neurosurgeon. As a result, the length of DSA sequences in the DIAS dataset ranges from 4 to 14 frames.
@@ III-B1 Full Annotation
Figure 5: Illustration of the proposed pipeline for DSA-sequence intracranial artery segmentation. (a) Vessel sequence segmentation network with (b) sequence feature extraction module; (c) WSL: scribble learning-based segmentation composed of scribble supervision and consistency regularization; (d) SSL: random patch-based self-training framework. (WSC: Weight Sharing Convolution, SF: Sequence Fusion, GMP: Global Max Pooling, WSD: Weight Sharing Downsampling, DA: Data augmentation, WA: Weak Augmentation, SA: Strong Augmentation)
Manual annotation of IA in DSA sequences requires taking into account all the image information in the sequence. Fig. 1 shows an example of intracranial vessels in five consecutive frames of the arterial phase in the sagittal view. During this period, the contrast agent is gradually dispersed from the internal carotid artery (ICA) to cerebral vessels such as the anterior cerebral artery (ACA), middle cerebral artery (MCA), and posterior cerebral artery (PCA) with blood flow. For a single image, the contrast agent is visible only in a portion of the IA. For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+4, it filled the ACA, MCA, and PCA. Annotating the IA of each image in the sequence and then merging the annotation mask can maximize the utilization of the sequence’s vessel information to obtain accurate annotations. The thinness of vessels, particularly in later-stage arterial images, renders the commonly used strategy of outlining their borders in medical image annotation unfeasible. Thus, we have to annotate each vessel pixel by pixel, which is a labor-intensive and time-consuming procedure. Fig. 3 shows that our proposal for a compromise solution, which reduces workload while maintaining annotation accuracy:
  • The arterial-phase DSA sequences were divided into two phases from the middle: the anterior phase and the later phase. The contrast agent during the anterior phase predominantly fills the ICA as opposed to the smaller vessels such as the ACA, MCA, and PCA arteries, which are filled by the contrast agent in the later phase.
@@
  • The annotations of thick and thin vessels are merged using a pixel-level OR operation (1 for vessel pixel and 0 for background) to create the ultimate vessel mask.
We randomly selected 60 sequences from DIAS for the purpose of annotating blood vessels, using the patient ID as the criterion for selection. The remaining 60 sequences were used for research on IA semi-supervised segmentation. The DSA sequence was initially annotated by two MD students of surgery, followed by corrections from an associate physician with 10 years of experience, and finally confirmed and double-checked by a chief physician with 15 years of experience.

### 17 Dec 2023 v3

# DIAS: A Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

@@ I Introduction
Cerebrovascular diseases substantially contribute to global mortality and long-term disability [1]. Recent decades have witnessed a decline in mortality from intracranial artery (IA) diseases like intracranial arterial stenosis (ICAS), middle cerebral artery occlusion (MCAO), and intracranial aneurysm, thanks to advancements in medical and imaging technology [2]. Among these, Digital Subtraction Angiography (DSA), an imaging modality, visualizes the flow of a contrast agent through vessels by capturing a sequence of images over a duration typically spanning 3 to 15 seconds, with a sampling rate between 3 to 5 frames per second [3]. It effectively removes bony structures by subtracting pre-contrast images, thereby enhancing the visibility of vessels filled with the contrast agent. Owing to its inherently superior spatial and temporal resolution, DSA can accurately reveal lesion details in instances where computed tomography angiography (CTA) and magnetic resonance angiography (MRA) may not offer a definitive diagnosis [2]. Thus, DSA is widely recognized as the gold standard for investigating lesion angioarchitecture, deciphering the dynamics of arterial blood supply, and guiding endovascular intervention [4].
@@
Conventionally, neuroradiologists analyze DSA images visually for diagnosis, clinical decision-making, and postoperative evaluation, which is inherently time-consuming, labor-intensive, and subjective. Automated vessel segmentation in DSA promises to alleviate these challenges by accurately outlining vascular morphology. Currently, DSA-based cerebrovascular image segmentation finds utility in various downstream clinical applications, including automated thrombolysis in cerebral infarction scoring [5], 3D reconstruction of vessel [6], and image guidance of interventional surgery [7]. Despite its critical importance, cerebrovascular segmentation in DSA has not been extensively researched. A key factor contributing to this is the lack of a dedicated DSA dataset specifically designed for developing and evaluating segmentation techniques. However, the creation of such a dataset is particularly challenging due to the unique characteristics of DSA imaging. Primarily, in the DSA sequence, the intravascular contrast agent is projected from a fixed angle, revealing only a portion of the intracranial arteries within the contrast medium in each image, as illustrated in Fig. 1. Annotating vessels in DSA necessitates neuroradiologists to meticulously examine each image frame, a process that is both labor-intensive and costly [8]. Moreover, the intricate nature of cerebral vessels, characterized by overlapping structures, reduced contrast, and complex background noise, significantly increases the difficulty of accurate annotation.
@@
Currently, several segmentation approaches for DSA cerebral vessels have emerged [9, 10, 8]. However, these studies rely on in-house datasets, lacking a unified standard for evaluation. Most of these menthods [9, 10] focus on segmenting single-frame images, thereby ignoring DSA sequences’ inherent temporal attributes. As a result, single-frame based segmentation approaches may not fully represent the intricate vascular morphology of IA, potentially compromising their effectiveness in the diagnosis and treatment of cerebrovascular diseases [8]. Vessel segmentation in DSA is considered a dimensionality reduction task, where the input is 2D+Time images, and the output is a 2D segmentation map. This approach deviates from the traditional 2D and 3D vessel segmentation methods [11, 12, 13]. The key challenge lies in accurately capturing the spatiotemporal information present within the angiography for precise vessel segmentation.
@@
Figure 1: A representative sample of intracranial vessels in lateral view from DSA, where t to t+5 are consecutive frames of the arterial phase. The red arrow in the upper left image indicates segments C1-C3 of the intracranial internal carotid artery (ICA), while the lower left image represents the anterior cerebral artery (ACA) and the middle cerebral artery (MCA).
@@
In this work, we propose the DSA-sequence Intracranial Artery Segmentation dataset (DIAS), meticulously annotated at the pixel level by neurosurgeons. Acknowledging the complexity and labor-intensiveness of vessel annotation in DSA sequences, we’ve provided two types of weak annotations for training samples to support research in weakly-supervised IA segmentation. To comprehensively evaluate the efficacy of prevalent segmentation methods like fully-supervised, weakly-supervised, and semi-supervised segmentation on the DIAS dataset, we have established a benchmark comprising a vessel sequence segmentation network, a scribble-based learning segmentation framework, and a random patch-based self-training semi-supervised segmentation pipeline. Extensive experiments with the DIAS set robust baselines, potentially advancing future algorithmic developments in cerebral vessel segmentation.
@@ II-A Vessel Segmentation Datasets
Vessel segmentation, a prominent research field in computer-aided medicine, has seen significant advancements recently, largely due to technological progress and the availability of annotated datasets. We have summarized the several publicly available vessel segmentation datasets in Table I. A representative example is retinal vessel segmentation, which has many publicly available datasets, such as colour fundus datasets (DRIVE [14], CHASE_DB1 [15], STARE [16], and HRF [17]) and OCTA datasets (ROSE1 and ROSE2 [18]). These datasets consist of 20-117 2D images with spatial resolution ranging from 304$\times$304 to 3504$\times$2336. However, other vessel segmentation datasets are relatively scarce [8]. CHUAC [19] and DCA1 [20] are coronary angiography datasets with resolutions of 189$\times$189 and 300$\times$300, respectively. ASOCA [21] and VESSEL12 [22] are the 3D vessel segmentation datasets of coronary artery CTA and lung CT from the challenges, in which only training set data is available.
The proliferation of such datasets has catalyzed a wave of innovative research in the area of segmentation, with scholars endeavoring to devise algorithms that surpass current methodologies. Until now, the medical imaging community lacked publicly available DSA datasets with vessel annotations, with only limited studies using proprietary data. DCVessel [9] comprises 30 DSA arterial-phase images at a resolution of 700$\times$700, while HEART [23] comprises 1092 DSA coronary images at 256$\times$256 resolution. Su et al. [8] manually segmented arteries and veins on 97 DSA series, covering both anteroposterior and lateral views, each with a resolution of 1024$\times$1024 pixels. Our contribution, DIAS, offers a comprehensive IA segmentation dataset from DSA sequences, comprising 120 sequences (753 frames) at 800$\times$800 pixel resolution, spanning lengths of 4 to 14 frames. The DIAS can serve as a reliable benchmark for the development of algorithms on IA segmentation, hopefully inspiring further research in this arena.
@@ II-B Vessel Segmentation Methods
In the past few years, deep learning has revolutionized vessel segmentation, outperforming traditional approaches like hand-crafted features [24], filtering-based models [25], and statistical models [26]. Deep learning-based methods, particularly UNet [27] and its numerous variants, have become the primary methods for vessel segmentation. These networks typically employ an encoder-decoder architecture with popular components, such as multi-scale design [13, 9, 28], deep supervision [29, 9, 12], attention mechanisms [30, 13, 31], atrous convolution [32, 29, 12, 9], and improved skip connections [9, 12]. Moreover, several novel neural networks [33, 34] have also succeeded in vessel segmentation by leveraging generative adversarial networks (GAN) or graph neural networks (GNN). Beyond network architectures, some research has focused on optimizing training strategies [35, 36] and modifying loss functions [37].
@@
However, there have been few studies on DSA vessel segmentation. Meng et al. [9] proposed a CNN-based multiscale dense CNN segmentation network for IA segmentation in single-frame DSA, utilizing multiscale atrous convolution, enhanced dense blocks, and redesigned skip connections. Zhang et al. [23] proposed a boundary enhancement and feature denoising module to improve the extraction of boundary information in DSA coronary segmentation. These two approaches are tailored for vessel segmentation in individual DSA frames. For vessel sequence, Hao et al. proposed the sequential vessel segmentation network (SVS-Net) for vessel segmentation in X-ray coronary angiography, which incorporates 3D convolutional layers to extract rich temporal-spatial features and employs channel attention blocks to learn discriminative features [38]. Su et al. proposed a spatio-temporal U-Net (ST U-Net) architecture which integrates convolutional gated recurrent units (ConvGRU) in the contracting branch of U-Net for cerebral artery and vein segmentation in DSA sequences [8]. While both methods achieve vessel segmentation within sequences, the former is not adaptable to variable sequences, and the latter is challenging to extend to other models.
@@
Fully annotating vessel datasets is both time-consuming and costly. Consequently, annotation-efficient segmentation methods based on weakly-supervised and semi-supervised have attracted significant attention from researchers. Weakly-supervised segmentation (WSS) trains models with sparse annotations instead of full ones. The common methods involves online labeling [39] via consistency checks with the mean teacher or using pseudo-labels supervision [40] derived from model predictions. Semi-supervised segmentation (SSS) aims to explore tremendous unlabeled data with supervision from limited labeled data. Recently, self-training methods [41, 42] have dominated this field. Furthermore, methods employing consistency regularization strategies [41, 43, 44] improve the generalization ability by encouraging high similarity in predictions from two perturbed networks for the same input image. To our knowledge, there is no existing literature on semi-supervised or weakly-supervised vessel segmentation in DSA sequences. Building upon our created dataset, we have conducted in-depth research and proposed innovative methods in three distinct areas: network models, WSS, and SSS. These approaches are specifically tailored to leverage the unique features of the dataset, addressing each segmentation paradigm comprehensively.
@@ III-A Dataset Description
The DIAS dataset is derived from DSA images obtained during ischemic stroke treatments at Beijing Tiantan Hospital. This dataset encompasses cases of ICAS and MCAO. The ethical committee at Beijing Tiantan Hospital authorized the use of all the data presented in this study. All images were retrospectively collected between January 2019 and December 2021 and then anonymized to protect the patients’ privacy by removing all identifiable clinical treatment details. The original DSA sequences were captured by the Siemens angiography machine during the neurointerventional procedures and stored as DICOM files. It contains 1027 2D-DSA sequences from 416 patients in anteroposterior view or lateral view. In this study, we retained 120 sequences (753 images) after excluding those with significant motion artifacts or incomplete arterial phase, or any duplicates. A total of 120 sequences were derived from 60 patients, among whom 24 were female and 36 were male. The average age of these patients was 62.5 years, with an age range from 40 to 88 years old. The number of frames varies between sequences (6—46 frames), but the frame rate is fixed (4 FPS). The sequences were then extracted into individual images with a resolution of 800$\times$800 pixels using RadiAnt. We retained all the arterial-phase frames and removed the pre-contrast, capillary-phase, and venous-phase frames under the guidance of a neurosurgeon. As a result, the length of DSA sequences in the DIAS dataset ranges from 4 to 14 frames.
@@ III-B1 Full Annotation
Figure 2: Full annotated process of DIAS.
The manual annotation of IA in DSA sequences necessitates a thorough consideration of all image data within the sequence. Fig. 1 illustrates a typical representation of intracranial vessels across five successive frames during the arterial phase in a lateral view. During this phase, the contrast agent progressively spreads from the internal carotid artery (ICA) into cerebral vessels like the anterior cerebral artery (ACA) middle cerebral artery (MCA), carried by the blood flow. For a single image, the contrast agent is discernible in only a portion of the IA. For instance, the contrast agent at t filled the C1-C3 segment of the ICA, while at t+3, it filled the ACA and MCA. To annotate the IA effectively, each image in the sequence must be individually annotated, and then these annotations should be integrated to leverage the full vessel information available in the sequence, thereby ensuring precision in annotation. However, the challenge arises with the fine nature of vessels, especially in images from later arterial stages, which makes the standard approach of demarcating their borders impractical for medical image annotation. Thus, we have to annotate each vessel pixel by pixel, which is a labor-intensive and time-consuming procedure. In reality, the angiographic representations of vessels in adjacent frame images exhibit a high degree of similarity, indicating that annotating every pixel in each frame image involves a substantial amount of repetitive work. Fig. 2 presents our proposed solution, a pragmatic compromise that aims to reduce the workload without compromising the accuracy of annotations:
Figure 3: Full annotated representative samples of DIAS at anteroposterior view and lateral view.
  • [leftmargin=*,itemsep=-4pt]
  • The arterial-phase DSA sequences were divided into two phases from the middle: the anterior phase and the posterior phase. The contrast agent during the anterior phase predominantly fills the ICA as opposed to the smaller vessels such as the ACA and MCA arteries, which are filled by the contrast agent in the posterior phase.
@@
  • The annotations of thick and thin vessels are merged using a pixel-level OR operation (1 for vessel pixel and 0 for background) to create the ultimate groudtruth.
We randomly selected 60 sequences from the DIAS dataset for the purpose of annotating vessels, using patient IDs as the criterion for selection. The remaining 60 sequences were used for research on semi-supervised segmentation of IA. The DSA sequence was initially annotated by two MD students specializing in neurosurgery, followed by corrections from an associate physician with 10 years of experience, and finally confirmed and double-checked by a chief physician with 15 years of experience. Fig. 3 presents fully annotated examples of intracranial arteries in anteroposterior and lateral views.
Figure 5: Illustration of the proposed benchmark for DSA-sequence intracranial artery segmentation. (a) Vessel sequence segmentation network with (b) sequence feature extraction module; (c) WSS: scribble learning-based segmentation composed of scribble supervision and consistency regularization; (d) SSS: random patch-based self-training framework. (WSC: Weight Sharing Convolution, SF: Sequence Fusion, GMP: Global Max Pooling, WSD: Weight Sharing Downsampling, DA: Data augmentation, WA: Weak Augmentation, SA: Strong Augmentation)